40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 13015     Entry time: Thu May 25 19:27:29 2017     In reply to: 13010     Reply to this: 13017
Author: gautam 
Type: Update 
Category: General 
Subject: Coil driver board noises  

[Koji, Gautam]

Summary: 

  • Attachment #1 shows the measured/modeled noise of the coil driver board (labelled ITMX). 
  • Measurement was made with "TEST" input (which is what the DAC drives) is connected to ground via 50ohm terminator, and "BIAS" input grounded.
  • The model tells us to expect a noise of the order of 5nV/rtHz - this is comparable to (or below) the input noise of the SR785, and even the SR560. So this measurement only serves to place an upper bound on the coil driver board noise.
  • There is some excess noise below 40Hz, would be interesting to see if this disappears with swapping out thick-film resistors for thin film ones.
  • The LISO model says that the dominant contribution is from the voltage and input current noise of the two op-amps (LT1125) in the bias LP filter path. 
  • But if we can indeed realize this noise level of ~10-20nV/rtHz, we are already at the ~10^-17m/rtHz displacement noise for MICH at about 200Hz. I suspect there are other noises that will prevent us from realizing this performance in displacement noise.

Details:

This measurement has been troublesome - I was plagued by large 60Hz harmonics (see Attachment #1), the cause of which was unknown. I powered all electronics used in the measurement set up from the same power strip (one of the new surge-protecting ones Steve recently acquired for us), but these remained present. Yesterday, Koji helped me troubleshoot this issue. We did the various things, I try to put them here in the order we did them:

  1. Double check that all electronics were indeed being powered from the same power strip - OK, but harmonics remained present.
  2. Tried using a different DC power supply - no effect.
  3. Checked the signal with an oscilloscope - got no additional insight.
  4. I was using a DB25 breakout board + pomona minigrabbers to measure the output signal and pipe it to the SR785. Koji suggested using twisted ribbon wire + soldered BNC connector (recycled from some used ones lying around the lab). The idea was to minimize stray radiation pickup. We also disconnected the WiFi extender and GPIB box from the analyzer and also disconnected these from the power - this finaly had the desired effect, the large harmonics vanished. 

Today, I tried to repeat the measurement, with the newly made twisted ribbon cable, but the large 60Hz harmonics were back. Then I realized we had also disconnected the WiFi extender and GPIB box yesterday.

Turns out that connecting the Prologix box to the SR785 (even with no power) is the culprit! Disconnecting the Prologix box makes these harmonics go away. I was using the box labelled "Santuzza.martian" (192.168.113.109), but I double-checked with the box labelled "vanna.martian" (192.168.113.105, also a different DC power supply adapter for the box), the effect is the same. I checked various combinations like 

  • GPIB box connected but not powered
  • GPIB box connected with no network cable

but it looks like connecting the GPIB box to the analyzer is what causes the problem. This was reproducible on both SR785s in the lab. So to make this measurement, I had to do things the painful way - acquire the spectrum by manually pushing buttons with the GPIB box disconnected, then re-connect the box and download the data using SRmeasure --getdata. I don't fully understand what is going on, especially since if the input connector is directly terminated using a 50ohm BNC terminator, there are no harmonics, regardless of whether the GPIB box is connected or not. But it is worth keeping this problem in mind for future low-noise measurements. My elog searches did not reveal past reports of similar problems, has anyone seen something like this before?

It also looks like my previous measurement of the de-whitening board noises was plagued by the same problem (I took all those spectra with the GPIB boxes connected). I will repeat this measurement.

Next steps:

At the meeting this week, it was decided that

  • All AD797s would be removed from de-whitening boards and also coil-driver boards (as they are unused).
  • Thick film resistors with the most dominant noise contributions to be replaced with thin-film ones.
  • Gain of 3 on de-whitening board to be changed to gain of 1.

I also think it would be a good idea to up the 100-ohm resistors in the bias path on the ITM coil driver boards to 1kohm wire-wound. Since the dominant noise on the coil-driver boards is from the voltage noise of the Op-Amps in the bias path, this would definitely be an improvement. Looking at the current values of the bias MEDM sliders, a 10x increase in the resistance for ITMX will not be possible (the yaw bias is ~-1.5V), but perhaps we can go for a 4x increase?

The plan is to then re-install the boards, and see if we can 

  1. Turn on the whitening successfully (I checked with an extender board that the switching of the whitening stages works - turning OFF the "simDW" filter in the coil driver filter banks enables the analog de-whitening).
  2. Relize the promised improvement in MICH displacement noise with the existing whitening configuration.

We can then take a call on how much to up the series resistance in the DAC signal path. 

Now that I have figured out the cause of the harmonics, I will also try and measure the combined electronics noise of de-whitening board + coil driver board and compare it to the model.

Quote:
  • The last piece (?) in this puzzle is the coil driver noise - this needs to be modeled and measured.

 

Attachment 1: coilDriverNoises.pdf  143 kB  | Hide | Hide all
coilDriverNoises.pdf
ELOG V3.1.3-